# Ensemble Learning for Large Language Models in Text and Code Generation: A Survey

### v1

@@ 3.1.3 Mixture of Experts
Architecture-level ensemble methods are used in transformers as expert networks. These methods help model the distributions of multidomain data for different tasks. Some of these tasks include text generation, mathematics, code generation, and natural language understanding (NLU) [21, 6, 64].
Given input token $x$ and $K$ expert networks, the output next-token $y$ of an expert layer is typically determined by the weighted sum of the outputs for each expert network $E_{i}(x)$, where the weights are given by the output of a gating network (router) $G(x)_{i}$, as shown in Figure 7. This relationship can be expressed mathematically as
@@
which shows that this model can be considered a mixture of experts (MoEs).
MoEs combine different expert distributions from various domain datasets using a gating mechanism. This approach allows MoEs to create more optimal distributions than individual experts [21]. They scale model sizes more effectively without a significant increase in computational complexity [85].
In the equation 12, the elements of $G(x)_{i}$ represent the top $n$ experts, where $n>1$. This selection ensures that the gating functions receive meaningful gradients [12]. The elements in $E_{i}(x)$ represent the $n$ expert sub-blocks, which may use different sets of weights, such as those in the SwiGLU architecture [86].
As the number of experts per token, $n$, increases, the model’s parameter count rises while keeping computational costs constant. This is because the computation in MoEs is largely sparse [21]. Alternatively, MoEs can be adapted to Switch Routing, where only one expert is selected from a reduced batch size of parallel experts [6]. This $n=1$ routing strategy simplifies implementation and reduces routing computation and communication costs.
@@ 3.1.4 Reward Ensemble
Architecture-level ensembling can be applied to reward modeling in reinforcement learning by using multiple reward models (RMs), known as a reward ensemble. This section discusses the reward ensemble within the context of reinforcement learning from human feedback (RLHF).
This approach was first introduced to address reward hacking, also known as overoptimization [22]. Reward hacking occurs when a policy gains high rewards while ignoring true human preferences [87]. The typical structure of the reward ensemble in the RLHF pipeline is shown in Figure 8.
In this model, the learned RMs aim to approximate human preferences, represented by the policy $\pi$. The policy $\pi$ is optimized in relation to the policy model and the RMs. This can be done using LLMs with different initialization seeds [22] or through low-rank adaptation [48] while supervised fine-tuning (SFT) or pretraining [50].
@@
where $d:=\sqrt{D_{KL}(\pi\ ||\ \pi_{init})}$. When using Mean Optimization method [88], the ensemble reward function of $R$ given $K$ LLMs can be written as
@@
where $R_{i}$ is the gold RM score given the prompt $q$ and its response $a$ sampled from $i$-th RM. In addition to this method, we can test various optimization techniques based on performance criteria. Examples include Worst-Case Optimization [88] and Uncertainty-Weighted Optimization [89].
Since computing RM ensembling is costly, a multi-head RM approach has been suggested. This method saves memory and time when training LLMs by using shared encoders with separate linear heads [49].
We can also extend the RM ensemble to include weight-averaged reward models using (5) and fine-tune multiple RMs with different hyperparameters. This strategy enhances scalability, improves robustness against label corruption, and increases reliability during distribution shifts while mitigating reward hacking [47].
@@ 3.2.1 Output Ensemble
  • Scoring model:
    A scoring model produces meaningful numeric values by either classifying data with an ML model or generating results with a language model. These models are commonly used in ranking and various text generation tasks, such as outputs ranking, instruction-following [53], QA reasoning [55], and image caption generation [79].
    For instance, a random forest (RF) classifier can be trained to give confidence scores for answers generated by LLMs. This helps in choosing the most confident response [55].
Additionally, scoring models are important for routing tasks.
@@ 4 Text and Code Generation using LLM Ensemble Methods
We have seen success in using LLM ensembles for various language understanding tasks. We believe this approach can also be applied to real-world text and code generation tasks. Our goal is to expand the use of LLM ensemble models in these areas while overcoming the limitations associated with single-LLM approaches mentioned earlier.
@@
LLM ensemble learning is a relatively new research topic. Until recently, its application to text and code generation had not been thoroughly explored. As the techniques have developed, several articles have been published on related topics such as instruction following [53, 23, 64, 58, 57, 59], reasoning [9, 6, 55], MMLU [61], and code generation [21, 24].
We selected these articles based on their performance results and potential for wider applications. The findings show that LLM ensembling techniques enhance the naturalness, accuracy, and quality of generated outputs. We categorize these LLM ensemble approaches into four methods based on their definitions in Section 3. This section will discuss the model structure, training methods, subjective evaluation, and benefits of these approaches.
4.1 Code Generation Using MoE Ensemble Models
In this approach, the MoE ensemble technique is used for input-to-expert routing in language generation. Input-to-expert routing identifies which experts to use based on the input tokens. It builds on traditional methods, such as gating networks in RNN-based machine translation [12] and NN-based vowel discrimination [98].
One example of this method is transformer-based code generation using MoEs. This approach improves on conventional LLM training [21]. It achieves two key improvements: First, it uses only 13 billion active parameters per token, outperforming larger models like Llama2, which uses 70 billion parameters per token. Second, the MoE layer in Figure 7 replaces the FFN sub-block in the transformer block, allowing for better learning of language features. To speed up execution on single GPUs, Megablocks [99] casts the FFN operations of the MoE layer as large sparse matrix multiplications. As for its training and implementation, remind Section 3.1.3.
We conducted subjective evaluations to demonstrate the effectiveness of this approach [21]. The results are summarized in Table 1, which shows the pass rates from a code functional correctness test on three models. For instance, 60.7% of the code generated by the Mixtral model passed the unit tests, outperforming the LLaMA model. The LLaMA 2 70B and GPT-3.5 models were built using transformers with 70 billion and 175 billion parameters, respectively. In contrast, the Mixtral model uses a transformer with 8 experts and 47 billion parameters, effectively modeling sparse mixture of expert networks at each layer of a decoder-only model. The table shows that Mixtral, with fewer parameters, achieved significantly better functional correctness than a single LLM with many more parameters.
This approach is not limited to code generation; it can also be applied to text generation tasks. See Table 4 for a list of applications areas of LLM ensemble models. For example, it can be used for instruction-following with the MoE transformer [64] and for common sense reasoning with Switch transformers [6]. Similar to the code generation method, input-to-expert routing is determined by the weighted outputs of the mixture experts in a gating network based on the input tokens. MoEs are then used to model the expert function between source and target tokens for each expert. The subjective evaluation results confirm the effectiveness of these approaches when using natural language tokens as features. Notably, for Switch transformers, a time-to-quality threshold score improved from 131.1 to 62.8 with the optimized architecture [6].
Table 1: The subjective pass rate (pass@1) among code generated using the LLaMA 2 70B, GPT-3.5, and Mixtral 8x7B models [21]
@@ 4.1 Code Generation Using MoE Ensemble Models
Dataset LLaMA 2 70B GPT-3.5 Mixtral 8x7B
MBPP 49.8% 52.2% 60.7%
HumanEval 29.3% - 40.2%
   Mixtral with fewer parameters achieved significantly better functional correctness than the use of a single LLM with a lot more parameters.
4.2 Text Generation Using Output Ensemble Model
This approach uses an output ensemble model to improve text generation by combining a ranker and a fuser. For instance, a post-ranking fusion framework with two separate language models has been proposed [53]. In this framework, the ranker assesses linguistic quality, while the fuser merges the top-ranked candidate outputs.
First, the input prompt $x$ is combined with the responses from $K$ LLMs using separator tokens to create the input data. Then, the generative model is trained to produce a final response $y^{*}$ by fusing the top-ranked responses (Figure 12).
Figure 12: Post-ranking fusion framework using output ensemble applied to an instruction-following task. It consists of a ranker and a fuser in 2 steps [53].
Specifically, the ranker’s input features include the instruction and a set of $N$ LLM output responses in English, formatted as a JavaScript array. The fuser’s input consists of the top $K$ responses from $N$ LLMs, which it uses to generate the final output. Each training sample is represented by a language feature vector, which includes multiple frames of text embedding (e.g., a 768-dimensional vector of floating-point numbers) and quality scores. A single transformer learns conversational text features, allowing to model the linguistic quality score between the instruction and response features. The ranker and fuser are trained separately in this framework.
To handle the different tasks of score prediction (creating a ranked list of texts) and text generation (producing the fused text), the approach employs a cross-attention transformer for the ranker and a fine-tuned T5 model [71] for the fuser. The transformer includes a multi-layer perceptron that takes encoded chat texts as input and candidates in a single-head layer. Training for the transformer ranker starts with supervised learning on the instruction dataset. Here, a multi-layer perceptron is trained using chat text features as observations. Transformer features are derived from the text embedding data, and various scoring functions, such as BERTScore [100] and BARTScore [91], are used to optimize quality scores. This is followed by supervised fine-tuning of the T5 fuser, where the input and $K$ output texts are concatenated using separator tokens to generate an enhanced output. This step models the relationship between the input instruction and each candidate response using a sequence-to-sequence prediction approach.
This training method has two main advantages over using a single LLM for generation: 1) Each component is flexible and can be adapted for different applications. For example, it can be extended for code generation by using a listwise ranker to improve output. Instead of ranking candidate responses against the input instruction, as in instruction-following tasks, a code generation task can train a listwise ranker [101] on public code datasets [102] to rank multiple code snippets by runtime, potentially reducing the runtime of the fused code. 2) The strong output from a closed-source LLM, like ChatGPT, can be improved through post-ranking fusion. This differs from architecture-level ensemble generation, which only use open-source LLMs. Since a powerful LLM can give a solid baseline, functional correctness of the fused code can be ensured.
Table 2 presents the natural language generation (NLG) scores for three instruction-finetuned LLMs: the T5 baseline (Flan-T5), the LLaMA2 baseline (Alpaca), and the output ensemble model using the proposed post-ranking fusion approach (llm-blender) [53]. Comparing llm-blender with Alpaca and Flan-T5 shows that the proposed method outperforms traditional single LLM baselines in modeling and generating conversational texts. The quality improvement from the baselines to llm-blender indicates that the strengths of multiple LLMs are effectively captured when diverse responses are combined.
Table 2: The subjective NLG scores among response texts generated using the llm-blender, Alpaca, and Flan-T5 models [53]
@@ 4.2 Text Generation Using Output Ensemble Model
Model BERTScore BARTScore BLEURT
llm-blender 79.09 -3.02 -0.17
Alpaca 71.46 -3.57 -0.53
Flan-T5 64.92 -4.57 -1.23
   llm-blender ensembling open-source LLMs outperforms the conventional single LLM baselines by improving conversational text quality.
4.3 Results Aggregation Using Output Ensemble Models
This approach, like the previous one, generates response texts based on input instructions using a training-free output ensemble model. However, instead of fine-tuning a language model to combine the top-$K$ LLM responses, this method prompts an LLM to synthesize responses generated from pre-selected LLMs.
A Mixture-of-Agent (MoA) approach has been proposed for text generation [23]. The MoA consists of multiple layers of LLM agents. Each layer updates the outputs from the previous layer to enhance performance. Figure 13 shows the structural model representation for a four-layer MoA. The top three layers of the MoA form a feed-forward structure that integrates responses in a left-to-right direction using the Aggregate-and-Synthesize prompt (Figure 14).
Figure 13: MoA structure with 4 layers with 3 LLM agents in each layer [23].
Figure 14: Aggregate-and-Synthesize prompt to integrate multi-response [23].
Each LLM first produces a response text to provide more context and diverse perspectives. Similar to the MoE-based approach discussed in Section 4.1, the MoA combines the roles of the gating network and expert networks, making it a model-level MoE. During synthesis, responses are refined iteratively by re-aggregating them at each layer. The responses from the last layer are fed to the aggregator LLM, which uses the Aggregate-and-Synthesize prompt to generate a single high-quality output.
This output ensemble method could also enhance code generation quality by addressing coding problems using the MoA structure and prompt augmentation techniques discussed in Section 3.3. For example, we can prompt programming questions to generate solutions in Python, C++, Java, JavaScript, and Go and test their functional correctness using HumanEval dataset [1]. A subjective instruction-following accuracy test compared the MoA-based model’s performance with the latest LLM [23]. The results are shown in Table 3. In this experiment, the latest LLM ranked first on the leaderboard as of July 22, 2024. AlpacaEval 2.0 and MT-Bench served as benchmarks for instruction-following tasks, evaluating models based on a nuanced understanding of correctness, relevance, and efficiency. The table shows that the MoA-based ensemble model with deeper layers achieved better quality than the latest single LLM according to human-validated accuracy scores.
Other approaches of output-ensemble-based QA reasoning can be found [55]. One alternative [55] combined output ensemble and prompt ensemble methods. It used CoT or RA prompts as inputs for output-ensemble-based QA reasoning. This QA prediction did not require prior knowledge of the question’s reasoning type; it only used a scoring classifier, making it more computationally cost-effective than MoA.
The LLM ensemble approach with output ensemble models has also been applied to other tasks, such as clinical text error detection [103], disease diagnoses [56, 104], and image retrieval [79]. A prompt-based results aggregation method was proposed for clinical text error detection [103]. In this method, subtasks of error detection and correction were all asked in a single prompt, aggregating results from the latest closed-source LLMs. A majority-vote method [56] and a frequency-based weighted method [104] have also been used in this prompt-based aggregation. In one study [56], JSON was utilized in CoT-augmented prompts to model nested responses and graph-like structures.
Table 3: The subjective instruction-following accuracy scores among response texts generated using MoA and MoA-Lite correspond to the 6 proposers with 3 layers and with 2 layer respectively, and GPT-4o mini [23]
@@ 4.3 Results Aggregation Using Output Ensemble Models
Model AlpacaEval2.0 LC Win. MT-Bench Avg.
MoA 65.1 9.25
MoA-Lite 59.3 9.18
GPT-4o mini (05/13) 57.5 9.19

### v2

@@ 3.1.3 Mixture of Experts
Architecture-level ensemble methods, known as Mixture of Experts (MoE), are employed in transformers to model multidomain data distributions for tasks such as text generation, mathematics, code generation, and natural language understanding (NLU) [21, 6, 64].
For an input token $x$ and $K$ expert networks $E_{i}(x)$, the output $y$ of an MoE layer is the weighted sum of expert outputs, with weights determined by a gating network (router) $G(x)i$ (see Figure 7). This is expressed as:
@@
MoE combines expert distributions from diverse datasets via a gating mechanism, creating more optimal distributions than individual experts [21]. This allows for effective model scaling without significantly increasing computational complexity [88].
In Equation 12, $G(x)_{i}$ represents the top $n$ experts ($n>1$), ensuring meaningful gradients for the gating functions [12]. $E_{i}(x)$ denotes the $n$ expert sub-blocks, potentially with distinct weights (e.g., SwiGLU architecture [89]).
Increasing the number of experts per token ($n$) raises the model’s parameter count with constant computational costs, due to sparse computation in MoE [21]. Alternatively, Switch Routing selects a single expert ($n=1$) from a reduced batch of parallel experts, simplifying implementation and lowering routing and communication costs [6].
@@ 3.1.4 Reward Ensemble
Architecture-level ensembling using multiple reward models (RMs), termed a reward ensemble, can be applied to reward modeling in reinforcement learning. This section discusses reward ensembles within reinforcement learning from human feedback (RLHF).
This approach was introduced to address reward hacking (overoptimization), where a policy achieves high rewards while disregarding true human preferences [22, 90]. The typical structure of a reward ensemble in the RLHF pipeline is depicted in Figure 8.
In this model, learned RMs approximate human preferences, represented by the policy $\pi$, which is optimized relative to the policy model and RMs. This optimization can involve LLMs with different initialization seeds [22] or low-rank adaptation during supervised fine-tuning (SFT) or pretraining [48, 50].
@@
where $d:=\sqrt{D_{KL}(\pi\ ||\ \pi_{init})}$. Using the Mean Optimization method [91], the ensemble reward function $R$ for $K$ LLMs is:
@@
where $R_{i}(q,a)$ is the gold RM score for prompt $q$ and its response $a$ sampled from the $i$-th RM. Besides Mean Optimization, other techniques can be evaluated based on performance criteria, such as Worst-Case Optimization [91] and Uncertainty-Weighted Optimization [92].
Due to the computational cost of RM ensembling, a multi-head RM approach, which uses shared encoders with separate linear heads, has been proposed to save memory and time during LLM training [49].
The RM ensemble can be extended to include weight-averaged reward models (using Equation 5) and by fine-tuning multiple RMs with varying hyperparameters. This enhances scalability, improves robustness against label corruption, and increases reliability during distribution shifts, while also mitigating reward hacking [47].
@@ 3.2.1 Output Ensemble
Scoring model. A scoring model produces meaningful numeric values by either classifying data with an ML model or generating results with a language model. These models are commonly used in ranking and various text generation tasks, such as outputs ranking, instruction-following [53], QA reasoning [55], and image caption generation [82].
For instance, a random forest (RF) classifier can be trained to give confidence scores for answers generated by LLMs. This helps in choosing the most confident response [55]. Additionally, scoring models are important for routing tasks.
@@ 4 Text and Code Generation using LLM Ensemble Methods
Given the success of LLM ensembles in various language understanding tasks, we believe this approach can also benefit real-world text and code generation. Our goal is to expand the application of LLM ensemble models in these areas, addressing the limitations of single-LLM approaches.
@@
LLM ensemble learning is an emerging research area. Its application to text and code generation has only recently been explored, with several publications covering related topics such as instruction following [53, 23, 32], math reasoning [29, 43, 20, 44, 9], MMLU [21, 60, 61], conversation [51, 48], and code generation [31, 32, 20, 44, 101, 23, 21].
We selected articles based on: (1) significance within each main ensemble category (Section 3), (2) provision of clear metrics and quantifiable results, and (3) representation of recent (2023-2024) and seminal works. This ensures coverage of established techniques and promising emerging methods. Our analysis indicates that LLM ensembling consistently enhances the naturalness, accuracy, and quality of generated text and code. Following our categorization framework, we examine 2-3 representative papers from each main ensemble approach, discussing their performance, benchmark results, computational requirements, and strengths/weaknesses. This focused selection allows for an in-depth analysis of impactful ensemble strategies and a comprehensive overview of the field’s evolution.
4.1 Evaluation Metrics and Datasets
To assess these ensemble methods, we first detail the standard metrics and datasets for text and code generation.
For text generation, common metrics include:
  • Accuracy: Correctness of responses.
  • Win Rate: Comparative preference over other models.
  • BLEU: N-gram overlap with references, e.g., for fluency.
  • Alignment Score: Adherence to instructions/user intent.
  • Human Evaluation Score: Subjective assessment of coherence, relevance, and quality.
Key datasets include MMLU (multitask understanding) [102], GSM8K (math reasoning) [103], AlpacaEval (instruction following) [4], and MT-Bench (multi-turn conversation) [104].
For coding, metrics focus on correctness and efficiency:
  • Pass@k: Probability of a correct solution in $k$ attempts.
  • CodeBLEU: Syntactic and semantic similarity for code.
  • Execution Time & CPU Usage: Code efficiency.
Prominent datasets are HumanEval [1] and MBPP (Mostly Basic Python Problems) [105] for testing functional correctness.
4.2 Performance Analysis by Method
Using this evaluation framework, we now analyze the performance, resources, and trade-offs of seven key LLM ensemble methods.
4.2.1 Weight Merging
Recent weight merging methods efficiently combine specialized language models without requiring additional training or extensive resources. We examine results from two papers: [32], which evaluates primary approaches including Task Arithmetic [29], TIES-Merging [31], and DARE [32] (Table 1), and [43], which advances the field using genetic algorithms to achieve state-of-the-art results (Table 2).
Key Findings. Weight merging approaches demonstrate remarkable effectiveness across diverse tasks. DARE [32] eliminates up to 99% of parameter differences without performance degradation, while different techniques show complementary strengths: Task Arithmetic [29] excels in math (66.34% on GSM8K), TIES [31] in code generation (37.80% on HumanEval), and TIES+DARE balances both capabilities. Though all merged approaches still trail GPT-4’s 85.3% on conversation tasks. Meanwhile, Evolutionary Optimization [43] uses genetic algorithms to discover optimal combinations by evolving architectures and merging weights, enabling three weak 7B models to collectively outperform GPT-3.5 on Japanese math reasoning (52.0% vs. 50.4%).
Resource Requirements. Both methods feature minimal computational demands compared to fine-tuning. DARE [32] operates on CPUs alone, while Task Arithmetic [29] and TIES Merging [31] require no training, functioning purely through weight manipulation. Evolutionary Optimization [43] needs only a single GPU for optimization, not training. Neither approach increases parameter count or requires additional training data, making them significantly more accessible than conventional model development techniques.
Trade-offs. Weight merging methods offer remarkable computational efficiency, enabling researchers with limited resources to combine specialized capabilities without training. DARE excels in parameter efficiency through sparsification, while Evolutionary Optimization automates merging discovery, outperforming GPT-3.5 with smaller models. However, these approaches are limited by source model capabilities and show task-dependent variations in performance. Despite these constraints, weight merging democratizes advanced model development while reducing computational costs.
Table 1: Performance comparison of primary weight merging approaches on text and code generation tasks [32]
@@ 4.2.1 Weight Merging
Model Type HumanEval AlpacaEval GSM8K
WizardLM-13B 1 36.59% 67.20% 2.20%
WizardMath-13B 2 64.22%
Task Arithmetic 1 + 2 28.66% 67.04% 66.34%
TIES Merging 1 + 2 37.80% 68.63% 15.77%
TIES + DARE 1 + 2 36.59% 68.70% 36.16%
GPT-4 (03/14) 85.3%
  Task Arithmetic [29] excels in math, TIES [31] in code, and TIES+DARE [32] balances both, though all trail GPT-4 on AlpacaEval.
Table 2: Performance comparison of Evolutionary-Optimization-based weight merging against source models and GPT-3.5 [43]
@@
Model Type JP-LMEH MGSM-JA
WizardMath-7B 3 60.1 18.4%
Shisa Gamma 7B 4 66.1 9.6%
Abel 7B 002 5 56.5 30.0%
EO 3 + 4 + 5 70.5 52.0%
GPT-3.5 50.4%
  EO with weak open-source models outperformed GPT-3.5, a commercial model, on Japanese datasets, showing a strong generalization ability.
4.2.2 Knowledge Fusion
Knowledge fusion combines complementary knowledge from language models with diverse architectures through lightweight continual training. Unlike weight merging (requiring architectural similarity) or aggregation methods (demanding parallel deployment), this approach supports fusion across structurally different models. Introduced in FuseLLM [20] and extended in FusionChat [44], it shows promising results across benchmarks (Table 3).
Key Findings. FUSIONCHAT-7B [44] effectively leverages strengths from three architecturally diverse models (NH2-Mixtral-8x7B, NH2-Solar-10.7B, OpenChat-3.5-7B). Unlike traditional knowledge distillation, the fusion model surpasses all constituent models in reasoning (8.00) and coding (6.15), achieving competitive results against closed-source GPT-3.5 despite its modest 7B parameter size.
Resource Requirements. Knowledge fusion employs a two-stage process: knowledge transfer via lightweight fine-tuning followed by parameter-space merging using variation-ratio-based weights [44]. With no size constraints on target models relative to sources, it requires modest computational resources, enabling creation of powerful composite models from diverse architectures without extensive pretraining datasets or prohibitive costs.
Trade-offs. While knowledge fusion creates systems exceeding individual source model performance, challenges remain in specialized areas where GPT-4 maintains advantages (8.55 vs 6.15 in coding). Recent advances in knowledge distillation techniques [106] suggest promising pathways to further enhance efficiency by addressing capacity gaps through dynamic distribution interpolation. Despite being bounded by the collective knowledge of constituent models, this approach offers a promising direction for democratizing advanced language capabilities without relying on raw scaling.
Table 3: Performance comparison of knowledge fusion approach on text and code generation tasks of mt-bench [44]
@@ 4.2.2 Knowledge Fusion
Model Type Writing Reasoning Coding
NH2-Mixtral-8x7B 1 9.70 6.65 6.00
NH2-Solar-10.7B 2 9.50 7.35 4.00
OpenChat-3.5-7B 3 9.00 7.75 5.50
FUSIONCHAT-7B VARM 1 + 2 + 3 9.20 8.00 6.15
GPT-3.5 (March) 9.20 5.65 6.90
  FUSIONCHAT [44] performs comparable to a closed-source GPT-3.5.
4.2.3 Mixture-of-Experts
Mixture-of-Experts (MoE) architectures dynamically route inputs through specialized neural subnetworks, enabling efficient scaling of model capacity without proportional computational costs. Unlike traditional dense transformers, MoE models selectively activate only relevant experts for each input token. Recent implementations like Mixtral [21] and OLMOE [101] demonstrate the effectiveness of this approach across diverse tasks (Table 4).
Key Findings. Mixtral 8x7B [21] achieves superior performance with significantly fewer active parameters, outperforming larger models like LLaMA 2 70B across multiple benchmarks. Despite using only 13B active parameters per token (from a total of 47B), it demonstrates stronger capabilities in MMLU (70.6%) and code generation (MBPP: 60.7%) compared to denser models. Similarly, OLMOE-1B-7B [101] shows remarkable efficiency, achieving state-of-the-art performance in code generation (HumanEval: 63.7%) while using just 1B active parameters per token.
Resource Requirements. MoE architectures employ a router network at each layer to dynamically select relevant experts, typically activating only a small subset (e.g., 2 out of 8 experts in Mixtral) per token. This sparse activation pattern enables efficient inference despite large total parameter counts. Training requires specialized infrastructure to handle expert parallelism, but inference costs remain competitive due to the sparse computation pattern.
Trade-offs. While MoE models demonstrate superior parameter efficiency and performance across many tasks, they introduce additional complexity in routing mechanisms and expert load balancing. The approach shows particular promise in specialized tasks like code generation, where OLMOE matches or exceeds GPT-3.5’s performance (63.7% vs 62.8% on HumanEval) with far fewer parameters. However, the need for specialized training infrastructure and careful expert design represents a notable implementation challenge compared to dense model.
Table 4: MoE Performance on text and code generation
[…]
Model MMLU MBPP HumanEval
LLaMA 2 70B 69.9% 49.8% 29.3%
GPT-3.5-Turbo 70.0% 52.2% 62.80%
Mixtral 8x7B [21] 70.6% 60.7% 40.2%
OLMOE-1B-7B [101] 56.3% 63.7%
  MoEs [21, 101] with fewer parameters achieved significantly better functional correctness than the use of a single LLM with a lot more parameters.
4.2.4 Reward Ensemble
Reward ensemble methods enhance reinforcement learning from human feedback (RLHF) by aggregating multiple reward models to improve prediction accuracy and uncertainty estimation. Recent work in efficient reward ensembles [51] and diverse LoRA approaches [48] demonstrates significant alignment improvements across benchmarks (Table 5).
Key Findings. LoRA-based reward ensembles [48] outperformed both single reward models and simple ensemble averaging across metrics. They increased alignment scores from below 4.16 to 4.35 on MT-Bench with Llama-7B and improved accuracy from 0.685 to 0.710 on the Anthropic Helpful dataset. These improvements were achieved without proportional computational increases, effectively addressing the overoptimization problem common in RLHF.
Resource Requirements. These approaches maintain efficiency through lightweight techniques like linear-layer ensembles and LoRA-based methods. Diverse LoRA ensembles [48] maximize parameter efficiency by focusing adaptations on critical model components while encouraging diversity through nuclear norm maximization, requiring substantially fewer resources than multiple full reward models.
Trade-offs. While reward ensembles improve alignment, they introduce integration challenges for real-world applications. Limited evaluation on standardized benchmarks complicates adoption decisions, and production deployment requires complex orchestration of multiple reward models. Despite these challenges, the approach shows strong potential for resource-constrained environments, as reinforcement learning generally delivers superior performance when properly tuned, particularly if future work addresses evaluation gaps with comprehensive benchmarking against industry-standard metrics.
Table 5: reward ensemble performance on text generation
[…]
Base Model Dataset Metric Single RM Ens. LoRA
Llama-7B [51] MT-Bench Alignment Score < 4.16 4.35
Llama2-7B [48] Anthropic Helpful Accuracy 0.685 0.710
  LoRA-based reward ensemble outperforms single-RM-based models.
4.2.5 Output Ensemble
Output ensemble techniques combine responses from multiple language models to produce higher-quality outputs than individual models alone. Unlike weight merging or reward ensembles, output ensembles operate at inference time by aggregating multiple outputs. The Mixture-of-Agents (MoA) approach [23] demonstrates how layered agent architectures leverage collective model strengths without architectural modifications. The HumanEval results in Table 6 were obtained through our experiments combining four models (Llama-3.3-70B-Instruct-Turbo, Mixtral-8x7B-Instruct-v0.1, Qwen2.5-Coder-32B-Instruct, and DeepSeek-R1-Distill-Llama-70B-free) with Qwen2.5-Coder-32B-Instruct as the aggregator using Together API.
Key Findings. MoA [23] substantially outperforms GPT-4o mini across multiple domains. On AlpacaEval2.0, it achieves a 65.1% win rate versus GPT-4o mini’s 57.5%. Similarly, it shows superior performance on alignment (MT-Bench: 9.25 vs 9.19) and code generation (HumanEval: 92.0% vs 86.5%). These improvements emerge from MoA’s layered architecture, where each agent refines responses by considering outputs from previous layers before a final aggregator synthesizes a coherent response.
Resource Requirements. Output ensemble approaches require no model modification or retraining. MoA implements a model-level Mixture-of-Experts architecture where distinct LLMs function through prompt-based interactions. This architecture supports heterogeneous models regardless of size, architecture, or access level. As demonstrated in our experiments, combinations of open-source LLMs can match or exceed the performance of strong closed-source commercial models at potentially lower total cost, offering an economically viable alternative to expensive single-model solutions.
Trade-offs. While output ensembles improve performance across benchmarks, they introduce additional latency due to multiple model calls. The approach has proven versatile in applications from clinical text error detection [107] and disease diagnosis [56, 108] to image retrieval [82] and instruction-following via a post-ranking fusion like LLM-Blender [53]. Despite increased inference complexity, output ensembles remain attractive for production environments where model updates are frequent, offering a practical path to high-quality language capabilities without expensive retraining.
Table 6: output ensemble performance on text and code generation
[…]
Model AlpacaEval2.0 MT-Bench HumanEval
GPT-4o-mini (05/13) 57.5% 9.19 86.5%
MoA 65.1% 9.25 92.0%